# AlphaViT: A Flexible Game-Playing AI for Multiple Games and Variable Board Sizes

### v2

@@ 3.1 Architectures of DNNs
AlphaViT’s DNN has a significant drawback: the size of the policy vector is fixed by the input size of the transformer encoder layer. To address this issue, we propose two improved DNN architectures for AlphaViD and AlphaVDA. Although their DNNs share many similarities with AlphaViT’s DNN, they employ both a transformer encoder and a transformer decoder to calculate the value and move probability.
@@
In AlphaViD and AlphaVDA, the input board is linearly embedded using a convolutional layer and fed into a transformer encoder, similar to AlphaViT. The input embedding sequence is defined as follows:
${\mathbf{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{N_{p}}_{p}\bm{\mathrm{E}}],$ (6)
@@
where the sequence consists of a value token, a game token, and patch embeddings. The embedding sequence in AlphaViD and AlphaVDA differs from that in AlphaViT. The estimated value is obtained from the value head that processes the output embedding corresponding to the value embedding from the last layer of the transformer encoder. In DNNs of AlphaViD and AlphaVDA, the encoder output is only used directly for value estimation, indicating no need for the embedding size to match the action-space size.
@@
The architecture of AlphaViD’s DNN is shown at the top of Fig. 3. The DNN estimates the move probability using the transformer decoder and MLPp. The input embeddings for the transformer decoder are derived from the outputs of the transformer encoder corresponding to the patch embeddings, which are further processed through a fully connected layer as follows:
${\mathbf{y}}^{\prime}_{0}=\mathrm{MLP}([\bm{z}^{2}_{L};...;\bm{z}^{N_{e}+1}_{L}]),{\mathbf{y}}^{\prime}_{0}\in R^{N_{d}\times D_{d}},$ (7)
@@
where $N_{d}$ is the number of embeddings, and $D_{d}$ is the embedding size. Since the embedding sequence size must match the input size of the transformer decoder, ${\mathbf{y}}^{\prime}_{0}$ is interpolated to ${\mathbf{y}}_{0}\in\mathbb{R}^{N_{a}\times D}$, where $N_{a}$ is the action space size. This interpolation provides flexibility to adjust the action space size depending on the game type and board size. Similar to the original transformer, the transformer decoder receives ${\mathbf{y}}_{0}$ and the output of the transformer encoder ${\mathbf{z}}_{L}$. Finally, MLPp calculates the move probabilities from the output of the last layer of the transformer decoder ${\mathbf{y}}_{L}$. The probability of action $a_{n}$ is given by:
$p(a_{n})=\mathrm{Sigmoid}(\mathrm{MLP}_{p}(\mathrm{LN}({\mathbf{y}}_{L})))(n),$ (8)
where $a_{n}$ is the $n$th action.
@@
The architecture of AlphaVDA’s DNN is shown at the bottom of Fig. 3. The DNN has a similar architecture to AlphaViD’s DNN but uses learnable embeddings ${\mathbf{y}}^{\prime}_{0}$ as input to the transformer decoder. While the sizes of these learnable embeddings are fixed, they are interpolated to ${\mathbf{y}}_{0}$ to match the action space size, ensuring compatibility and adaptability across different game configurations.
@@
To decide on a move, AlphaViD and AlphaVDA use Monte Carlo Tree Search (MCTS) with Upper Confidence Bound (UCT), using the value and move probabilities predicted by their DNNs. The MCTS algorithm implemented in these agents is identical to that of AlphaZero, as detailed in Appendix 1. The hyperparameters for AlphaViD and AlphaVDA are listed in Appendix 3.
@@ 4.1 Games
This study evaluates the performance of AlphaViT and AlphaViD across six games: Connect4, Connect4 5x4, Gomoku, Gomoku 6x6, Othello, and Othello 6x6. These games are two-player, deterministic, zero-sum games with perfect information. Connect4, published by Milton, is a connection game played on a $7\times 6$ board. The players take turns dropping discs onto the board. A player wins by forming a straight line of four discs horizontally, vertically, or diagonally. Connect4 5x4 is a variant of Connect4 with a $5\times 4$ board. Gomoku is a connection game in which players place stones on a board to form a line of five stones in a row, either horizontally, vertically, or diagonally. This study uses a $9\times 9$ board for Gomoku and a $6\times 6$ board for Gomoku 6x6. Othello (Reversi) is a two-player strategy game played on an $8\times 8$ board. In Othello, the disc is white on one side and black on the other. Players take turns placing a disc with their assigned color facing up. During a game, discs of the opponent’s color are flipped to the current player’s color if they are in a straight line and bounded by the disc just placed and another disc of the current player’s color. Othello 6x6 is played on a $6\times 6$ board in this study.
@@ 4.2 Opponents
This study evaluates the performance of AlphaViT, AlphaViD, and AlphaVDA using five different AI methods: AlphaZero, two variants of Monte Carlo Tree Search (MCTS) labeled MCTS100 and MCTS400, Minimax, and Random. AlphaZero is trained using the method described in Appendix 1. The MCTS methods (MCTS100 and MCTS400) were implemented using different numbers of simulations (100 and 400, respectively). The details of MCTSs are provided in Appendix 4. In these MCTS methods, the child nodes are expanded at the fifth visit to a node. Minimax selects a move using the minimax algorithm based on the evaluation table described in Appendix 5. Random selects moves uniformly at random from valid moves.
@@ 4.3 Software
AlphaViT, AlphaViD, AlphaVDA, the opponents, and the board games were implemented in Python, using NumPy for linear algebra operations and PyTorch for deep neural networks. The source code is available on GitHub at https://github.com/KazuhisaFujita/AlphaViT for reproducibility and further extension of this work.
@@ 5 Results
The results section presents the performance and characteristics of the proposed AI agents: AlphaViT, AlphaViD, and AlphaVDA. These agents are trained on different games (Connect4, Gomoku, and Othello) with two board sizes (small and large). The architectures of these agents vary primarily according to the number of encoder layers, which influences their capability to learn from the game environment. Table 1 shows the number of parameters for each agent configuration. AlphaViT, AlphaViD, and AlphaVDA are tested with different numbers of encoder layers, denoted as ‘L’ followed by a number (e.g., L1, L4, L5, L8). The number of parameters ranges from 11.2 million to 19.9 million, increasing with the number of encoder layers. For comparison, the AlphaZero agent, which serves as the baseline, has 7.1 million parameters.
Each AI agent is trained on specific games with different board sizes. Table 2 categorizes the agents based on the games on which they are trained and the board sizes used during training. The first group includes agents trained on a single game with a large board, denoted as LB. The second group consists of agents trained on a single game with a small board, denoted as SB. The third group comprises agents trained on multiple games, including Connect4, Gomoku, and Othello, with large boards, denoted as Multi. The agents in the third group are trained simultaneously on the three games and can play these three games using a single DNN. In other words, they do not specialize in a specific game. This diversity in training regimes allows us to evaluate the adaptability and generalization capabilities of the agents across different game domains.
@@
AI agent num of encoder layers num of parameters
@@ 5.1 Baseline Elo ratings
Table 3 lists the Elo ratings of various AI agents for different games and board sizes. Elo ratings provide a standard measure of relative performance in two-player games and enable systematic comparison of the performance of each AI agent. The evaluated agents include variations of the proposed methods trained on a single game with a large board (LB), a single game with a small board (SB), and multiple games with large boards (Multi). These are compared with other AI agents, including AlphaZero, Monte Carlo Tree Search (MCTS) with different numbers of simulations, Minimax, and a Random agent. The proposed agents and AlphaZero are trained over 1000 iterations. The Elo ratings of all agents are initialized to 1500 and calculated through 50 round-robin tournaments between the agents. The details of the Elo rating calculation are provided in Appendix 6.
AlphaViT L4 LB achieves strong performance, achieving Elo ratings of 1824 in Connect4, 1835 in Gomoku, and 2125 in Othello. However, for Connect4 and Gomoku, AlphaZero outperforms AlphaViT. Interestingly, despite not being trained on small-board games, AlphaViT L4 LB performs as well as or better than Minimax and MCTS100 in Connect4 5x4, Gomoku 6x6, and Othello 6x6. Furthermore, AlphaViT L4 LB exhibits competitive performance for Othello with AlphaViD L1 SB, AlphaVDA L1 SB, and AlphaZero. This suggests that AlphaViT may efficiently utilize the knowledge obtained from training on large-board games.
AlphaViT L4 SB shows Elo ratings of 1732 in Connect4 5x4, 1764 in Gomoku 6x6, and 1969 in Othello 6x6. In these games, AlphaViT L4 SB performs as well as or better than AlphaZero. For Gomoku and Othello, it outperforms Minimax and MCTS even though it is not trained on large board games. This result suggests that AlphaViT L4 SB may effectively utilize knowledge from training on small-board games for large-board games.
AlphaViT L4 Multi, trained simultaneously on Connect4, Gomoku, and Othello, shows strong results in these games, achieving Elo ratings of 2003 in Connect4, 1969 in Gomoku, and 1939 in Othello. Although it is weaker than AlphaZero in Connect4 and Gomoku, AlphaViT L4 Multi outperforms AlphaViT L4 LB (trained on Gomoku only) in Gomoku. Furthermore, AlphaViT L4 Multi exhibits comparable performance to MCTS400 in Gomoku 6x6.
AlphaViD L1 LB performs slightly worse than AlphaViT in the games with the large boards. For Gomoku 6x6 and Othello 6x6, AlphaViD L1 LB outperforms Minimax and MCTS100, suggesting that it effectively transfers knowledge gained from training on large boards. AlphaViD L1 SB achieves Elo ratings of 1742 in Connect4 5x4, 1785 in Gomoku 6x6, and 1852 in Othello 6x6, with performance comparable to AlphaZero and AlphaViT. AlphaViD L1 Multi shows performance close to AlphaViT in Connect4 but is weaker in Gomoku and Othello.
AlphaVDA L1 LB performs worse than AlphaViT for the games with large boards. AlphaVDA L1 SB shows Elo ratings of 1770 in Connect4 5x4, 1871 in Gomoku 6x6, and 1833 in Othello 6x6, with performance comparable to AlphaZero, AlphaViT, and AlphaViD. AlphaVDA L1 Multi is on par with or slightly weaker than AlphaViD L1 Multi.
The Elo ratings presented in Tab. 3 serve as a baseline for the subsequent experiments described in the following sections.
Table 3: Elo Ratings of AI Agents Across Different Games and Board Sizes
Game Connect4 Connect4 5x4 Gomoku Gomoku 6x6 Othello Othello 6x6
AlphaViT L4 LB 1824 1477 1835 1572 2125 1820
AlphaViD L1 LB 1755 1472 1724 1532 1856 1581
AlphaVDA L1 LB 1757 1506 1594 1244 1745 1336
AlphaViT L4 SB 1169 1732 1564 1764 1487 1969
AlphaViD L1 SB 1187 1742 1503 1785 1128 1852
AlphaVDA L1 SB 1213 1770 973.5 1785 1131 1833
AlphaViT L4 Multi 2003 1374 1969 1654 1939 1136
AlphaViD L1 Multi 1743 1295 1538 1392 1695 1379
AlphaVDA L1 Multi 1720 1321 1524 1194 1565 953.2
AlphaZero 2058 1767 2279 1769 1889 1850
Minimax 1033 1337 1484 1403 1127 1385
MCTS100 1260 1512 1153 1414 1192 1372
MCTS400 1550 1581 1169 1675 1365 1642
Random 725.8 1057 697.0 838.6 741.5 883.2

### v3

@@ 3.1 Architectures of DNNs
AlphaViT’s DNN has a significant drawback: the size of the policy vector is fixed by the transformer encoder’s input size. To address this issue, AlphaViD and AlphaVDA incorporate both a transformer encoder and a decoder for calculating value and move probabilities, as shown in Figure 3. In AlphaViD, the decoder receives the input derived from the output of the encoder. In contrast, AlphaVDA employs learnable action embeddings as inputs to the decoder. The final output of the encoder is used to compute the value, while the final output of the decoder is used to compute the move probabilities.
Figure 3: Architecture diagrams of AlphaViD (top) and AlphaVDA (bottom). The input board is divided into patches, embedded, and combined with value and game tokens, then processed by a transformer encoder. For policy prediction, encoder outputs are sent to a transformer decoder via interpolation. In AlphaViD, decoder inputs are generated from board-derived embeddings via a fully connected layer; in AlphaVDA, learnable action embeddings are used. MLP heads output the final value and move probabilities.
@@
In AlphaViD and AlphaVDA, the input board is linearly embedded using a convolutional layer and fed into a transformer encoder, similar to AlphaViT. However, unlike AlphaViT, no pass embedding is included in the encoder input. The input embedding sequence is defined in Eq. 6 as follows:
$\mathbf{z}_{0}=[{\mathbf{x}}_{\mathrm{value}};{\mathbf{x}}_{\mathrm{game}};\mathbf{z}^{\mathrm{pos}}_{\mathrm{patch}}],$ (6)
@@
where the sequence consists of a value embedding, a game embedding, and patch embeddings. The embedding sequence in AlphaViD and AlphaVDA differs from that in AlphaViT. The estimated value is obtained from the value head that processes the output embedding corresponding to the value embedding from the last layer of the transformer encoder. In DNNs of AlphaViD and AlphaVDA, the encoder output is only used directly for value estimation, indicating that the number of patch embeddings $N_{p}$ does not need to match the action-space size.
@@
The architecture of AlphaViD’s DNN is shown at the top of Figure 3. The DNN estimates the move probability using the transformer decoder and MLPp. The input embeddings for the transformer decoder are derived from the outputs of the transformer encoder corresponding to the patch embeddings, which are further processed through a fully connected layer. The initial embeddings for the decoder input are defined in Eq. 7 as follows:
${\mathbf{y}}^{\prime}_{0}=\mathrm{MLP}([\mathbf{z}^{2}_{L};\ldots;\mathbf{z}^{N_{p}+1}_{L}]),\quad{\mathbf{y}}^{\prime}_{0}\in\mathbb{R}^{N_{p}\times D_{d}},$ (7)
@@
where $D_{d}$ is the embedding size of the decoder. Since the embedding sequence size must match the input size of the transformer decoder, ${\mathbf{y}}^{\prime}_{0}$ is interpolated to ${\mathbf{y}}_{0}\in\mathbb{R}^{N_{a}\times D_{d}}$, where $N_{a}$ is the action space size. In this study, this step uses bilinear interpolation in the function torch.nn.functional.interpolate. This interpolation provides flexibility to adjust the action space size depending on the game type and board size. If an additional move such as a “pass” (e.g., in Othello) is required, we simply set $N_{a}=HW+1$ and resize ${\mathbf{y}}^{\prime}_{0}$ accordingly via this interpolation. The additional embedding at the last position is assigned to the additional move (e.g., the pass). Therefore, a dedicated pass embedding is not required for the encoder’s input. Similar to the original transformer, the transformer decoder receives ${\mathbf{y}}_{0}$ as the target sequence and $\mathbf{z}_{L}$ as the memory (encoder output). Finally, MLPp calculates the move probabilities from the output of the last layer of the transformer decoder ${\mathbf{y}}_{L}$. The probability of action $a_{i}$ is given by Eq. 8:
$p(a_{i}\mid s)=\mathrm{Softmax}(\mathrm{MLP}_{p}(\mathrm{LN}({\mathbf{y}}_{L})))_{i}.$ (8)
@@
The architecture of AlphaVDA’s DNN is shown at the bottom of Figure 3. The DNN is similar to AlphaViD’s but uses learnable embeddings ${\mathbf{y}}^{\prime}_{0}$ as the initial embeddings for the decoder input to the transformer decoder. While the length of these learnable embeddings is fixed, they are interpolated to the decoder input ${\mathbf{y}}_{0}$ to match the action space size, ensuring compatibility and adaptability across different game configurations.
@@
To decide on a move, AlphaViD and AlphaVDA employ MCTS with UCT, using the value and move probabilities predicted by their DNNs. The MCTS algorithm implemented in these agents is identical to that of AlphaZero, as detailed in Appendix A. The hyperparameters for AlphaViD and AlphaVDA are listed in Appendix C.
@@ 4.1 Games
This study evaluates our agents on six game variants (three different games, each played on two board sizes): Connect 4 (7×67$\times$ 67 × 6; “Connect 4”, and 5×45$\times$ 45 × 4; “Connect 4 5x4”), Gomoku ($\times 9$ 9; “Gomoku”, and $\times 6$ 6; “Gomoku 6x6”), and Othello ($\times 8$ 8; “Othello”, and $\times 6$ 6; “Othello 6x6”). These games are two-player, deterministic, zero-sum games with perfect information. Connect 4, published by Milton Bradley, is a connection game played on a 7×67$\times$ 67 × 6 board. The players take turns dropping discs onto the board. A player wins by forming a straight line of four discs horizontally, vertically, or diagonally. Connect 4 5x4 is a variant of Connect 4 with a 5×45$\times$ 45 × 4 board. Gomoku is a connection game in which players place stones on a board to form a straight line of five stones, either horizontally, vertically, or diagonally. This study uses a $\times 9$ 9 board for Gomoku and a $\times 6$ 6 board for Gomoku 6x6. Othello (Reversi) is a two-player strategy game played on an $\times 8$ 8 board. In Othello, the disc is white on one side and black on the other. Players take turns placing a disc with their assigned color facing up. During a game, discs of the opponent’s color are flipped to the current player’s color if they are in a straight line and bounded by the disc just placed and another disc of the current player’s color. Othello 6x6 is played on a $\times 6$ 6 board in this study. We deliberately start with Connect 4, Gomoku, and Othello because (i) each has natural board-size variants (5x4 – 9x9), (ii) they are commonly used for benchmarking, and (iii) training can be completed on commodity GPUs.
@@ 4.2 Opponents
This study evaluates the performance of AlphaViT, AlphaViD, and AlphaVDA using five different AI methods: AlphaZero, two variants of MCTS labeled MCTS100 and MCTS400, Minimax, and Random. AlphaZero is trained using the method described in Appendix A. The MCTS methods (MCTS100 and MCTS400) were run with 100 and 400 simulations, respectively.
[…]
The details of MCTSs are provided in Appendix D. In these MCTS methods, the child nodes are expanded at the fifth visit to a node. Minimax selects a move using the minimax algorithm based on the evaluation table described in Appendix E. Random selects moves uniformly at random from valid moves.
@@ 4.3 Software
AlphaViT, AlphaViD, AlphaVDA, the opponents, and the board games were implemented in Python, using NumPy for linear algebra operations and PyTorch for DNNs. The source code is available on GitHub at https://github.com/KazuhisaFujita/AlphaViT for reproducibility and further extension of this work.
4.4 Hardware
All experiments were conducted on multiple custom-built PCs. The agents used a variety of consumer CPUs, including AMD Ryzen 9 9900X and Intel Core i7-14700K, Core i9-13900K, and Core i9-12900K processors. The experimental GPUs were NVIDIA GeForce RTX 4060 Ti 16 GB, RTX 3060 12 GB, RTX 2070 Super, and RTX 2070. Each machine was equipped with 64 GB of RAM and two GPUs, and the experiments were run on Debian Linux. The model training and evaluation were distributed across multiple GPUs on one machine using data parallelization. No cloud-based or high-performance computing clusters were used, and all computations were performed using local custom-built hardware. For example, one of the machines used for training and evaluation was equipped with an Intel Core i9-13900K CPU, 64 GB of RAM, and two NVIDIA GeForce RTX 4060 Ti 16 GB GPUs.
@@ 5 Results
The results section presents the performance and characteristics of AlphaViT, AlphaViD, and AlphaVDA which were trained on different games (Connect 4, Gomoku, and Othello) with two board sizes (large and small). The main architectural difference between these agents lies in the number of encoder layers, which directly affects their learning capacity. Table 1 shows the number of parameters for each agent configuration. Each agent was tested with different numbers of encoder layers, denoted by ‘L’ followed by a number (e.g., L1, L4, L5, L8). The number of parameters ranges from 11.2 to 19.9 million, increasing with encoder depth. For comparison, the AlphaZero agent, which serves as the baseline agent, has 7.1 million parameters.
The primary architectural variable is the number of transformer encoder layers. Throughout this paper, we designate four layers for AlphaViT and one layer for AlphaViD and AlphaVDA (11$\approx$ 11≈ 11 million parameters) as the baseline configuration, which we also refer to as the shallower configuration in contrast to the deeper variants described later. By contrast, deeper encoders use exactly four additional layers (e.g., L8 for AlphaViT, L5 for AlphaViD/AlphaVDA, 20$\approx$ 20≈ 20 million parameters). Here, “deeper” is used only in this relative sense and does not imply a universal threshold on the number of layers.
Each AI agent was trained on specific games with different board sizes. Table 2 categorizes the agents based on the games on which they were trained and the board sizes used during training. The first group includes agents trained on a single game with a large board, denoted as LB. The second group consists of agents trained on a single game with a small board, denoted as SB. The third group comprises agents simultaneously trained on multiple games, including Connect 4, Gomoku, and Othello, with large boards, denoted as Multi. The agents in the third group were trained on the three games and can play these three games using a single DNN. In other words, they do not specialize in a specific game. This diversity of training settings allows us to evaluate the agents’ adaptability and generalization capabilities across different game domains.
@@
AI agent Number of encoder layers Number of parameters
@@ 5.1 Baseline Elo ratings
Tables 3, 4, and 5 present Elo ratings of various AI agents across different games and board sizes. Elo ratings provide a standard measure of relative performance in two-player games and enable systematic comparison across agents. We evaluated multiple variants of our proposed models, including agents trained on a single game with either large or small boards, and multiple games with large boards. Comparisons were conducted with other AI agents, including AlphaZero, MCTS with different numbers of simulations, Minimax, and a Random agent. Our proposed agents and AlphaZero underwent 1000 training iterations; each iteration consisted of self-play, augmentation, and an update step detailed in Appendix B. The Elo ratings of all agents were initialized to 1500 and were calculated through 50 round-robin tournaments between the agents. Each Elo rating is accompanied by a 95% confidence interval (95% CI); the calculation details are provided in Appendix F.
Note: Throughout this paper, we use the term strong to indicate benchmark-relative performance: an agent is considered strong if its Elo rating closely approaches or matches that of our AlphaZero (see Tables 3, 4, and 5). Thus, strong does not refer to an absolute numerical threshold but instead to AlphaZero-level performance.
Highlights across games.
Single-task AlphaViT L4 approaches AlphaZero’s performance across all evaluated games, remaining within approximately 270 Elo points of AlphaZero. Single-task AlphaViD L1 and AlphaVDA L1 achieve AlphaZero-level performance only on smaller boards (Connect 4 5x4 and Gomoku 6x6), yet they experience substantial drops in Elo ratings on larger boards. Multitask AlphaViT L4 (AlphaViT L4 Multi) demonstrates competitive performance on large boards and closely approaches AlphaZero (within 200 Elo points). Conversely, multitask AlphaViD L1 and AlphaVDA L1 variants exhibit significantly lower performance. Note that the multitask models were not trained on small board configurations and therefore underperform on this board size.
Connect 4 (Table 3).
For standard Connect 4, AlphaViT L4 Multi most closely approaches AlphaZero ($\Delta$ Elo 160$\approx$-160≈ - 160). AlphaViT L4 LB is moderately weaker ($\Delta$ Elo 260$\approx$-260≈ - 260), whereas AlphaViD L1 LB and AlphaVDA L1 LB lag behind significantly ($\Delta$ Elo <300<-300< - 300). On the smaller 5x4 board, all single-task variants (AlphaViT L4 SB, AlphaViD L1 SB, AlphaVDA L1 SB) are competitive with AlphaZero.
Gomoku (Table 4).
For the large board, only AlphaViTs (L4 LB and Multi) attain performance comparable to AlphaZero, with overlapping confidence intervals. AlphaViD and AlphaVDA variants (L1 LB and Multi) substantially lag behind ($\Delta$ Elo <300<-300< - 300). On the smaller board, SB variants from all architectures effectively approach AlphaZero-level performance.
Othello (shown in Table 5).
On the large board, AlphaViT L4 LB numerically surpasses AlphaZero ($\Delta$ Elo +20$\approx$+20≈ + 20), though overlapping confidence intervals prevent definitive conclusions. AlphaViD L1 LB and AlphaViT L4 Multi achieve performance near that of AlphaZero, with $\Delta$ Elo 100$\approx$-100≈ - 100. AlphaVDA L1 LB lags behind significantly ($\Delta$ Elo <300<-300< - 300). On the smaller board, AlphaViT L4 SB closely approaches AlphaZero ($\Delta$ Elo 70$\approx$-70≈ - 70). AlphaViD L1 SB lags behind ($\Delta$ Elo 230$\approx$-230≈ - 230), whereas AlphaVDA L1 SB also falls short ($\Delta$ Elo 180$\approx$-180≈ - 180), remaining weaker than AlphaZero. Interestingly, AlphaViT L4 LB, despite being trained only on the large board, achieves performance comparable to AlphaViD L1 SB and AlphaVDA L1 SB on the smaller board. This suggests that knowledge may transfer from larger to smaller boards.
Conclusion.
In summary, AlphaViT L4 and AlphaViT L4 Multi consistently achieve or closely approach AlphaZero-level strength across games, even without game-specific fine-tuning. In contrast, AlphaViD L1 and AlphaVDA L1 achieve AlphaZero-level performance only on smaller boards. The strong results for multitask agents (AlphaViT L4 Multi) suggest that multitask training will not significantly hinder performance. The Elo ratings presented in Table 3, 4, and 5 serve as a baseline for the subsequent experiments described in the following sections.
Table 3: Elo ratings of AI agents for Connect 4 variants
Agent Connect 4 Connect 4 5x4
Elo $\Delta$ Elo 95% CI Elo $\Delta$ Elo 95% CI
AlphaZero ${\bf 2114}$ [207320732073,215921592159] ${\bf 1769}$ [174717471747,179717971797]
AlphaViT L4 LB 184618461846 267.4-267.4- 267.4 [179917991799,190219021902] 151715171517 252-252- 252 [148014801480,155415541554]
AlphaViD L1 LB 173917391739 374.7-374.7- 374.7 [169616961696,178317831783] 146214621462 306.9-306.9- 306.9 [142614261426,149514951495]
AlphaVDA L1 LB 174617461746 367.3-367.3- 367.3 [170717071707,178817881788] 150715071507 262.1-262.1- 262.1 [147214721472,153915391539]
AlphaViT L4 SB 120212021202 911.6-911.6- 911.6 [115211521152,125212521252] ${\bf 1751}$ 18.07-18.07- 18.07 [172217221722,177717771777]
AlphaViD L1 SB 117711771177 936.4-936.4- 936.4 [113511351135,121412141214] ${\bf 1764}$ 4.512-4.512- 4.512 [174117411741,179317931793]
AlphaVDA L1 SB 125112511251 862.5-862.5- 862.5 [120112011201,130113011301] ${\bf 1789}$ 19.9819.9819.98 [176317631763,181818181818]
AlphaViT L4 Multi 195019501950 164.0-164.0- 164.0 [191719171917,198819881988] 140214021402 366.4-366.4- 366.4 [136413641364,143714371437]
AlphaViD L1 Multi 174517451745 369.2-369.2- 369.2 [170617061706,178517851785] 131713171317 451.4-451.4- 451.4 [127112711271,136113611361]
AlphaVDA L1 Multi 166916691669 444.7-444.7- 444.7 [162716271627,171017101710] 130513051305 463.3-463.3- 463.3 [126612661266,134313431343]
MCTS400 151615161516 836.1-836.1- 836.1 [146714671467,156015601560] 156315631563 205.1-205.1- 205.1 [153015301530,159915991599]
MCTS100 127812781278 598.2-598.2- 598.2 [123012301230,131613161316] 150215021502 266.9-266.9- 266.9 [146814681468,153315331533]
Minimax 101210121012 1102-1102- 1102 [970.6970.6970.6,105310531053] 131613161316 453-453- 453 [127712771277,134613461346]
Random 755.4755.4755.4 1358-1358- 1358 [722.5722.5722.5,788.9788.9788.9] 103810381038 730.4-730.4- 730.4 [992.4992.4992.4,108210821082]
Elo ratings of AI agents for Connect 4 variants. Bolded ratings indicate agents whose Elo rating is within 100 points of AlphaZero. Elo ratings were calculated from 50 round-robin tournaments beginning with an initial rating of 1500, including differences relative to AlphaZero ($\Delta$ Elo) and corresponding 95% confidence intervals.
Table 4: Elo ratings of AI agents for Gomoku variants
Agent Gomoku Gomoku 6x6
Elo $\Delta$ Elo 95% CI Elo $\Delta$ Elo 95% CI
AlphaZero ${\bf 2038}$ [198219821982,209120912091] ${\bf 1807}$ [178817881788,183018301830]
AlphaViT L4 LB ${\bf 1966}$ 72.4-72.4- 72.4 [192419241924,201220122012] 164516451645 162-162- 162 [161116111611,167816781678]
AlphaViD L1 LB 169816981698 340-340- 340 [165916591659,174417441744] 150915091509 298.3-298.3- 298.3 [147614761476,154615461546]
AlphaVDA L1 LB 156715671567 471.5-471.5- 471.5 [152115211521,161416141614] 126112611261 546.3-546.3- 546.3 [121912191219,129712971297]
AlphaViT L4 SB 155315531553 485.7-485.7- 485.7 [150015001500,160516051605] ${\bf 1764}$ 42.78-42.78- 42.78 [174517451745,178617861786]
AlphaViD L1 SB 151915191519 519.6-519.6- 519.6 [146614661466,156515651565] ${\bf 1779}$ 27.43-27.43- 27.43 [176017601760,179817981798]
AlphaVDA L1 SB 977.2977.2977.2 1061-1061- 1061 [943.8943.8943.8,101110111011] ${\bf 1771}$ 35.96-35.96- 35.96 [175117511751,179217921792]
AlphaViT L4 Multi ${\bf 2024}$ 13.93-13.93- 13.93 [198519851985,206920692069] 165616561656 150.8-150.8- 150.8 [162616261626,168816881688]
AlphaViD L1 Multi 153015301530 508.7-508.7- 508.7 [148414841484,157615761576] 137813781378 428.8-428.8- 428.8 [133913391339,141614161416]
AlphaVDA L1 Multi 157015701570 467.8-467.8- 467.8 [151715171517,162216221622] 117011701170 636.9-636.9- 636.9 [113011301130,120912091209]
MCTS400 116111611161 877.4-877.4- 877.4 [111611161116,120112011201] 168016801680 126.3-126.3- 126.3 [165116511651,170917091709]
MCTS100 122912291229 809.5-809.5- 809.5 [118011801180,127412741274] 138013801380 426.3-426.3- 426.3 [134313431343,141814181418]
Minimax 147214721472 566.4-566.4- 566.4 [142514251425,151815181518] 134813481348 458.6-458.6- 458.6 [130913091309,138413841384]
Random 696.8696.8696.8 1341-1341- 1341 [673.9673.9673.9,719.2719.2719.2] 851.9851.9851.9 954.9-954.9- 954.9 [820.8820.8820.8,884.9884.9884.9]
Elo ratings of AI agents for Gomoku variants. Bolded ratings indicate agents whose Elo rating is within 100 points of AlphaZero. Elo ratings were calculated from 50 round-robin tournaments beginning with an initial rating of 1500, including differences relative to AlphaZero ($\Delta$ Elo) and corresponding 95% confidence intervals.
Table 5: Elo ratings of AI agents for Othello variants
Agent Othello Othello 6x6
Elo $\Delta$ Elo 95% CI Elo $\Delta$ Elo 95% CI
AlphaZero ${\bf 1996}$ [195319531953,204820482048] ${\bf 2034}$ [198619861986,208120812081]
AlphaViT L4 LB ${\bf 2017}$ 20.5520.5520.55 [197319731973,206520652065] 181918191819 215.5-215.5- 215.5 [177617761776,186318631863]
AlphaViD L1 LB ${\bf 1896}$ 99.95-99.95- 99.95 [185418541854,194119411941] 158515851585 449.5-449.5- 449.5 [153915391539,163116311631]
AlphaVDA L1 LB 166916691669 327.2-327.2- 327.2 [162016201620,171817181718] 136813681368 666.4-666.4- 666.4 [132213221322,141514151415]
AlphaViT L4 SB 148214821482 513.9-513.9- 513.9 [143514351435,152815281528] ${\bf 1963}$ 71.16-71.16- 71.16 [191919191919,201320132013]
AlphaViD L1 SB 118411841184 812.4-812.4- 812.4 [113211321132,122812281228] 180318031803 231.6-231.6- 231.6 [175717571757,185418541854]
AlphaVDA L1 SB $111$ 885.3-885.3- 885.3 [106310631063,115911591159] 185518551855 179.6-179.6- 179.6 [179917991799,191119111911]
AlphaViT L4 Multi ${\bf 1910}$ 86.4-86.4- 86.4 [186518651865,195819581958] 109010901090 944.2-944.2- 944.2 [104610461046,113411341134]
AlphaViD L1 Multi 166816681668 328.7-328.7- 328.7 [162616261626,170917091709] 130813081308 726.4-726.4- 726.4 [126412641264,135413541354]
AlphaVDA L1 Multi 157815781578 418-418- 418 [152915291529,162616261626] 991.1991.1991.1 1043-1043- 1043 [938.4938.4938.4,103410341034]
MCTS400 137313731373 623.6-623.6- 623.6 [132913291329,141214121412] 157915791579 455-455- 455 [153515351535,162316231623]
MCTS100 119411941194 802.7-802.7- 802.7 [114811481148,124112411241] 136713671367 667.6-667.6- 667.6 [132413241324,141214121412]
Minimax 114011401140 855.9-855.9- 855.9 [108910891089,119011901190] 136313631363 670.9-670.9- 670.9 [131113111311,141114111411]
Random 782.8782.8782.8 1213-1213- 1213 [744.8744.8744.8,816.5816.5816.5] 874.2874.2874.2 1160-1160- 1160 [824824824,919919919]
Elo ratings of AI agents for Othello variants. Bolded ratings indicate agents whose Elo rating is within 100 points of AlphaZero. Elo ratings were calculated from 50 round-robin tournaments beginning with an initial rating of 1500, including differences relative to AlphaZero ($\Delta$ Elo) and corresponding 95% confidence intervals.